OpenStudy (yoongilife):

I am willing to work. I want to learn. I will post document as well as possible answers once the post becomes active.

OpenStudy (yoongilife):

Ok, so I know the standard algebra formula y = mx + b. This is of course is used when you are given the slope + two points on the graph/line. But, I have reason to believe that y = mx + b is the incorrect way to go about solving this problem. For instance, y = 7x + b --- 0 = 7 (4) + b --- 0 = 28 + b --- -28 – 0 = 28 -28 + b --- y = 7x – 28. This is obviously not an answer choice, but I would like to learn. I am willing to contribute to the problem. Possible answers: (4, 6) (0, 2) (6, 4) (2, 0)

OpenStudy (mayankdevnani):

first write the equation of line

OpenStudy (yoongilife):

y = 7x - 28?

OpenStudy (mayankdevnani):

\[\large \bf eq.of~line \rightarrow y-y_1=m(x-x_1)\] where, \[\large \bf (x_1,y_1)~are~co-ordinates~through~which~line~is~passing\]

OpenStudy (yoongilife):

y - 0 = m ( x - 4)?

OpenStudy (mayankdevnani):

correct !

OpenStudy (mayankdevnani):

substitute `m` also !

OpenStudy (yoongilife):

y - 0 = 2 ( x - 4)...

OpenStudy (mayankdevnani):

Good job !

OpenStudy (mayankdevnani):

now, all you have to do is satisfy points in the equation of line,if they satisfy the equation, then that point is your possible answer :)

OpenStudy (yoongilife):

So, we would technically 'plug' in the values, observing if they fulfilled the equation or not?

OpenStudy (mayankdevnani):

yup

OpenStudy (yoongilife):

Would it be ok, if I test one out using one of the answers?

OpenStudy (mayankdevnani):

why not

OpenStudy (yoongilife):

Quick question, would I distribute the 2 (distributive property) or work out the equation within the parentheses first and then multiple? Thank you.

OpenStudy (mayankdevnani):

its your choice

OpenStudy (mayankdevnani):

its better to work within the parentheses

OpenStudy (yoongilife):

Working with (4, 6) I get 6 = 2, which is not correct. Can I keep going done the line @mayankdevnani?

OpenStudy (mayankdevnani):

yupp

OpenStudy (yoongilife):

Correct answer: (6, 4).

OpenStudy (mayankdevnani):

Good job :)

OpenStudy (yoongilife):

Thank you for your help! Would it be ok if I asked you a fairly similar question @mayankdevnani? I understand if your not available.

OpenStudy (mayankdevnani):

you can post it(only few minutes i am available) Not from earth... lol

OpenStudy (yoongilife):

Ok @mayankdevnani. If you get around to it I appreciate it.

OpenStudy (mayankdevnani):

Consider two points :- \[\large \bf (x_1,y_1)~and~(x_2,y_2)\] \[\large \bf Slope=\frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (yoongilife):

-1 + 5 = m ( -1 - 2) -1 + 5 / -1 + 2

OpenStudy (yoongilife):

The top one is the equation of the line and the bottom one is y2 -y1 / x2 - x1. Please let me know if I'm wrong, I would be more than happy to learn...

OpenStudy (mayankdevnani):

your equation of line is wrong dude

OpenStudy (mayankdevnani):

where is variable x and y ??

OpenStudy (mayankdevnani):

and for determining slope, you don't need to write the equation of line

OpenStudy (yoongilife):

Ok, your right an equation of a line is NOT needed. In that case if we followed the y2-y1 / x2-x1 formula, and 'pluged' in the points given to us from the problem wouldn't it be -1 + 5 / -1 - 2 (accounting for the double negative rule)?

OpenStudy (mayankdevnani):

correct !!

OpenStudy (yoongilife):

In that case our answer should be 4/-3 as the slope, right?

OpenStudy (mayankdevnani):

yupp

OpenStudy (yoongilife):

Thanks again! I have one more problem just like the one above. I would work it out myself, but would like you to check it. Before you said you only had a few minutes so I understand if you can't. Thanks again...

OpenStudy (mayankdevnani):

your welcome :)

OpenStudy (mayankdevnani):

i have to go now !!

OpenStudy (yoongilife):

Ok. Thanks!

OpenStudy (mayankdevnani):

Nice to see you again :)